Mike is 5 years old. His brother Shawn is three times as old. When Shawn is 20, how old will Mike be?
Answer by TimothyLamb(4379)

mike = 5
shawn = 3*5 = 15
---
difference in age = 15 - 5 = 10
---
When Shawn is 20, Mike will be = 20 - 10 = 10
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---